### v5

@@ 3 Bilateral Trade War for Competitiveness
[…]
Furthermore, country $i$ would choose:
@@
as $i$’s best trade partner and:
@@
as $i$’s worst competitor. In Eqs. (13) or (14), $\pi_{i}$ has the largest increase or decrease, respectively, in percentage per one percentage rise of $P_{ji}$. However, $i$ is not his best partner’s best partner nor his worst competitor’s worst competitor, according to Corollary 3. Thus, countries $i$ and $j$ are not in a sports-like competition or zero-sum game unless they are evenly matched. For example, in the 2018 China-USA trade war, both the Trump and Biden administrations actively challenged China, whereas China passively defended itself.
Eq. (11) describes the side effects on all other countries due to country $i$’s change in $P_{ji}$. With the perturbation in $P_{ji}$, some countries gain, while others lose their competitiveness; the gain or loss may be more substantial than $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$. Although $\pi_{i}$ and $\pi_{j}$ always move in opposite directions, the movements are not of the same magnitude. One movement could be significant, while the other remains insignificant; it is also possible that both movements are insignificant. Third parties unevenly share the difference between these two movements, and their individual or aggregate shares may be significant.
The side effects may play a role in determining $\lambda_{ji}$ when $i$ is much more potent than $j$. The conflict seems resolved if the first-order condition $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{i}$. However, if $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{j}$ instead, then in theory, $i$ could extremely counteract and push $\lambda_{ji}$ near zero. In practice, however, $i$ is unlikely to exercise these extreme actions because its gain from $j$ is capped by $\pi_{j}P_{ji}$, as seen from Eq. (4), which could be negligible to $i$. Also, $\pi_{j}P_{ji}$ could be less than the loss of third parties, measured by Eq. (11). So, excessively non-cooperative counteractions could provoke outrage from these countries, and $j$ could ally with them to seek collective bargaining with $i$. It could also pursue an objective other than maximizing $\pi_{j}$, avoiding confrontation with $i$. For example, it can engage in country $i$’s economy by developing complementary industries and focusing on its comparative advantages. For centuries, many norms and standards have been established to tame unilateral trade bullies; today, the World Trade Organization (WTO) deals with trade rules between nations.
@@ 4 Nash Bargaining Solutions for Reprisal Coefficients $\Lambda$
When additional factors are considered, the condition $\lambda_{ji}>\pi_{j}/\pi_{i}$ alone is insufficient for country $i$ to initiate a conflict against country $j$. Among these factors, trade deficits are paramount. This section explores bilateral bargaining solutions for $\lambda_{ji}$ when trade balance is also an objective.
@@ 4.1 Trade Balance
For a given period, a trade surplus for country $i$ with country $j$ implies a trade deficit of the same magnitude for $j$ with $i$, and vice versa. A trade surplus generates job opportunities, reduces unemployment, and expands the economy of scale. It also enhances creditworthiness by enabling the country to pay down foreign debt. A surplus can result from competitive advantages in tradable goods and services. Opinions on trade deficits vary, though large deficits can undermine economic sustainability. In the short term, trade deficits help avoid shortages and mitigate issues like inflation and poverty. The implications of trade deficits also depend on their impact on national security and how they are financed. No nation can completely ignore trade deficits or competitiveness; a one-sided emphasis will not endure.
There are two benchmark choices for $\lambda_{ji}$ regarding net trade balances, $g_{j}/g_{i}$ and $P_{ij}/P_{ji}$, as stated in Theorem 3. For country $i$ to have a trade surplus with country $j$ due to the change $\Delta P_{ij}$, $g_{i}\Delta P_{ij}>g_{j}\Delta P_{ji}$. Thus, $\lambda_{ji}>g_{j}/g_{i}$ implies $i$’s trade surplus if $\Delta P_{ji}>0$. To sustain trade peace with $\Delta P_{ji}>0$, a large $\lambda_{ji}$ improves $i$’s trade balance but harms $\pi_{i}$. Excessive exports from $i$ to $j$ transfer power to $j$, reducing $\pi_{i}$. Therefore, $\lambda_{ji}$ should be neither too large nor too small. However, $\lambda_{ji}<g_{j}/g_{i}$ also indicates $i$’s surplus if $\Delta P_{ji}<0$. In this case, both $i$ and $j$ lose exports, though $i$ loses fewer. This could adversely affect future production, so most countries avoid mutually damaging stalemates. Generally, $\lambda_{ji}$ is determined by both players, but the sign and size of $\Delta P_{ji}$ are often decided by one, especially if it has overwhelming power.
@@
$\lambda_{ji}=g_{j}/g_{i}$ or $g_{j}/g_{i}=P_{ij}/P_{ji}$ maintains a zero net trade deficit and surplus between countries $i$ and $j$, at the instant or cumulative level, respectively.
To ease trade tension, we have studied two ways to determine $\lambda_{ji}$. On the one hand, if both sides aim for higher competitiveness, then $\lambda_{ji}=\pi_{j}/\pi_{i}$ is the steady-state solution for any minor change in $\Delta P_{ji}$. On the other hand, if they seek balanced trade, then $\lambda_{ji}=g_{j}/g_{i}$ eliminates any future trade deficits. Generally speaking, $\lambda_{ji}=\pi_{j}/\pi_{i}$ is for long-term rivalry due to $\pi$’s involvement in all countries and industries; $\lambda_{ji}=g_{j}/g_{i}$ could immediately resolve the trade deficit and alleviate related problems such as unemployment and national debt. According to Pozsar (2022), for example, the 2018 China-USA trade war began in this manner: China had grown exponentially over four decades by exporting less expensive goods overseas, whereas the USA lost much of its manufacturing industries but accumulated massive national debt through wars on terror, significant tax cuts, and the 2007 subprime mortgage crisis. Nevertheless, when China sought to build a global 5G network and produce cutting-edge computer chips, the USA attempted to block the move and urged its allies to join the effort. In this example, the focal point is competitiveness rather than trade balance. Therefore, the conflict is expected to last long, and $\pi_{j}/\pi_{i}$ is an appropriate choice for $\lambda_{ji}$.
We examine a scenario where a superpower country $i$ faces imminent challenges from an emerging economy $j$. According to the Matthew effect, it is likely that $\pi_{j}/\pi_{i}<g_{j}/g_{i}$. Based on Theorem 4, country $i$ cannot simultaneously have an increasing $\pi_{i}$ and a net trade surplus while growing its exports to country $j$. This situation occurs only when $\Delta P_{ij}<0$, causing both countries to lose comparative advantages. Figure II illustrates this trilemma, where up to two policy objectives are achievable for country $i$. For instance, if country $i$ adopts a position on edge $b$, it can overturn comparative advantage and earn a trade surplus but sacrifices competitiveness unless it absorbs competitiveness from other countries. In the early 2000s, bilateral exports between China and the USA increased as a percentage of their respective production. According to the trilemma, the USA had to either lose competitiveness, incur trade deficits, or both.
@@
Given $\pi_{j}/\pi_{i}<g_{j}/g_{i}$, country $i$ could not simultaneously achieve the three objectives in its bilateral trade with country $j$: increasing $\pi_{i}$; trade surplus; growing exports (i.e., $\Delta P_{ij}>0$).
Figure II: impossibility trilemma of $\pi_{i}\uparrow$, trade surplus, and $\Delta P_{ij}>0$.
Country $i$’s interests in increasing $\pi_{i}$, trade surplus, and $\Delta P_{ij}>0$ are not well aligned and do not completely contradict those of country $j$. In Figure II, country $i$’s optimal policy interests could lie at any point inside the triangle, whereas country $j$’s interests could lie at another point. A common ground is that $\Delta P_{ij}>0$ and $\Delta P_{ji}>0$ when both countries benefit from escalating economies of scale and diminishing marginal costs of production. According to the proof of Theorem 4, country $j$ could simultaneously achieve all three objectives listed in the theorem. This seems unfair to country $i$, but it is conditional on the Matthew effect, which is somewhat biased toward $\pi_{i}$. Therefore, Theorem 4 serves as a redress to the effect.
@@ 4.2 Nash Bargaining Solutions
This subsection examines a compromise solution between trade balance and competitiveness. In the trade system, some countries prioritize competitiveness while others aim to improve trade balance, either by reducing deficits or increasing surpluses. Their preferences influence the choice of $\lambda_{ji}$, which maintains $\pi_{i}$ and trade balance unchanged at $\lambda_{ji}=\pi_{j}/\pi_{i}$ and $\lambda_{ji}=g_{j}/g_{i}$, respectively. If countries are indifferent between competitiveness and trade balance, the midpoint appears to be a solution, i.e., $\lambda_{ji}=\left(g_{j}/g_{i}+\pi_{j}/\pi_{i}\right)/2$ and $\lambda_{ij}=\left(g_{i}/g_{j}+\pi_{i}/\pi_{j}\right)/2$. However, this midpoint solution does not satisfy the identity condition $\lambda_{ji}\lambda_{ij}=1$. Alternatively, when countries $i$ and $j$ have opposing priorities between enhancing competitiveness and improving trade balance, the Nash bargaining solution (1950) can be employed to find a cooperative solution that balances $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ while adhering to the condition $\lambda_{ji}\lambda_{ij}=1$.
Let
@@
which are either $\pi_{j}/\pi_{i}$ or $g_{j}/g_{i}$. By Corollary 2 and Theorem 3, without bargaining, $\lambda_{ji}$ can secure $p$ and $q$, and $\lambda_{ij}$ can secure $1/p$ and $1/q$ because:
@@
We consider two scenarios when countries $i$ and $j$ have opposite preferences on $p$ and $q$. In the first scenario, $j$ prefers $q$ to $p$, whereas $i$ prefers $1/p$ to $1/q$. However, $(q,1/p)$ is not a solution for $(\lambda_{ji},\lambda_{ij})$ since $q\times 1/p\neq 1$. By Eq. (15), $p$ and $1/q$ are their respective status quo points, obtained if one decides not to bargain with the other. In the second scenario of opposite preferences, $j$ prefers $p$ to $q$ while $i$ prefers $1/q$ to $1/p$.
Theorem 5 addresses the bargaining problems, and the solution satisfies the identity $\lambda_{ji}\lambda_{ij}=1$ by design. The solution is consistent across both preference scenarios, thus it remains effective when either country $i$ or $j$ is indifferent between greater competitiveness and improved trade balance. The solution remains unchanged under any positive affine transformation of their utility functions, according to a property of the bargaining solution. By taking the square root of $\pi_{j}/\pi_{i}$, $\lambda_{ji}^{*}$ mitigates the Matthew effect in $\pi$. Since $\lambda_{ji}^{*}$ is between $g_{j}/g_{i}$ and $\pi_{j}/\pi_{i}$, the mitigation bridges the gap between these two ratios.
@@
When countries $i$ and $j$ have opposite preferences on $p$ and $q$, the Nash bargaining solution is:
@@
Figure III illustrates the bargaining solution. In the left plot (first scenario), $j$ moves from left to right along the curve $\lambda_{ij}=1/\lambda_{ji}$ while $i$ moves from bottom to top along the same curve. When they meet at the solution point $(\lambda^{*}_{ji},\lambda^{*}_{ij})$, the area of the rectangle cornered by $(\lambda^{*}_{ji},\lambda^{*}_{ij})$ and the status quo point $(p,1/q)$ is maximized. The right plot (second scenario) illustrates this situation where $j$ moves from $q$ to $p$ and $i$ from $1/p$ to $1/q$ along the curve. They meet at the solution point $(\lambda^{*}_{ji},\lambda_{ij}^{*})$, which maximizes the area of the rectangle, cornered by $(\lambda^{*}_{ji},\lambda^{*}_{ij})$ and $(q,1/p)$.
Figure III: Nash bargaining solution for $\lambda_{ji}^{*}$ when $i$ and $j$ have opposite preferences.
The solutions $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$ rely on the choice of status quo points. The effective bargaining set is $[p,q]$ for $\lambda_{ji}$ and $[1/q,1/p]$ for $\lambda_{ij}$ because country $j$ can secure $p$ (or $q$) and $i$ can secure $1/q$ (or $1/p$) in the first (or second, respectively) scenario. If both players intensify or weaken their status quo positions symmetrically, we scale these two status quo points by a positive number $c\leq\sqrt{q/p}$. Then, $\lambda_{ji}$’s effective bargaining set extends from $[p,q]$ to $[cp,q/c]$ and $\lambda_{ij}$’s extends from $[1/q,1/p]$ to $[c/q,1/(cp)]$. However, the bargaining solution remains unchanged, according to Corollary 4. Letting $c\to 0$, the new effective bargaining sets become $(0,\infty)$ while the symmetry still bounds the solution of $\lambda_{ji}^{*}$ within $[p,q]$ and $\lambda_{ij}^{*}$ within $[1/q,1/p]$.
@@
For any positive $c\leq\sqrt{q/p}$, if we replace $p$ with $cp$ and $q$ with $q/c$, the bargaining solution remains $\lambda_{ji}=\lambda_{ji}^{*}$ and $\lambda_{ij}=\lambda_{ij}^{*}$.
Additionally, as Corollary 5 states, ${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ are also Nash bargaining solutions for $\lambda_{ji}$, as implied in Corollary 2 and Theorem 3. In summary, when countries $i$ and $j$’s preferences are known, we recommend using the following $\lambda_{ji}$:
$\lambda_{ji}=\left\{\begin{array}[]{ll}\frac{\pi_{j}}{\pi_{i}},&\mathrm{if\ both\ prefer\ higher\ competitiveness;}\\ \frac{g_{j}}{g_{i}},&\mathrm{if\ both\ prefer\ better\ trade\ balance;}\\ \sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}},&\mathrm{otherwise.}\end{array}\right.$ (16)
Otherwise, we suggest using $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ as a rule of thumb.
@@
${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ are Nash bargaining solutions for $\lambda_{ji}$.
Finally, country $j$ should avoid direct confrontation with a significantly larger-sized economy $i$, according to Corollary 6. This was likely China’s strategy and practice from 1980 to 2010 when dealing with greater economic powers; Deng Xiaoping set its diplomatic policy to “keep a low profile and take no lead” (e.g., Yan, 2014), and trade policy is often used as a tool of foreign policy. Consequently, an even distribution of economic, political, and military capabilities between contending countries is likely to increase the chance of war, and peace is preserved best when there is an imbalance of national capabilities between them (e.g., Organski, 1968).
@@
Country $j$ should avoid homogeneous preferences with a significantly larger-sized country $i$.

### v6

@@ 3.2 Identification of Competitors and Collaborators
When Eq. (12) exceeds a positive threshold $\theta$ (e.g., $\theta=.1$) or falls below $\theta$, the derivative in Eq. (9) is considered significantly positive or negative, respectively. For instance, if Eq. (12) equals .2.2, then a $1\%$ increase in $P_{ji}$ would approximately lead to a $2\%$ increase in $\pi_{i}$. Therefore, a significantly negative (or positive) derivative in Eq. (9) serves as a necessary — but not sufficient — condition for country $i$ to initiate a trade war (or form an economic alliance) with country $j$, even if Eq. (10) remains insignificant.
Furthermore, country $i$ would select
@@
as its best trade partner and
@@
as its worst competitor. In Eqs. (13) and (14), $\pi_{i}$ experiences the largest increase or decrease, respectively, in percentage terms for each percentage rise in $P_{ji}$. However, according to Corollary 3, country $i$ is not necessarily its best partner’s best partner, nor its worst competitor’s worst competitor. This implies that international trade dynamics are not akin to sports-like competitions or zero-sum games — unless the countries involved are evenly matched. For example, during the 2018 China-USA trade war, both the Trump and Biden administrations actively challenged China, while China largely adopted a defensive posture.
Eq. (11) describes the spillover effects on other countries resulting from country $i$’s change in $P_{ji}$. When $P_{ji}$ is perturbed, some countries gain competitiveness while others lose it; these gains or losses can be more substantial than $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$. Although $\pi_{i}$ and $\pi_{j}$ always move in opposite directions, their magnitude differ: one change may be significant while the other remains negligible, or both may be minor. Third countries share the difference between these movements unevenly, and their individual or collective impacts can be substantial.
These spillover effects may influence the determination of $\lambda_{ji}$ when country $i$ is significantly more powerful than country $j$. The conflict appears resolved if the first-order condition $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{i}$. However, if this condition instead maximizes $\pi_{j}$, then, in theory, $i$ could aggressively counteract. In practice, such extreme actions are unlikely because $i$’s gain from $j$ is capped by $\pi_{j}P_{ji}$ (as shown in Eq. (4)), which may be negligible for $i$. Moreover, $\pi_{j}P_{ji}$ could be smaller than the losses incurred by third parties, as measured by Eq. (11). Excessively non-cooperative behavior could provoke backlash from these countries, potentially leading $j$ to form alliances and pursue collective bargaining against $i$.
Alternatively, $j$ might adopt objectives beyond maximizing $\pi_{j}$, such as avoiding direct confrontation with $i$. For instance, it could deepen economic engagement with $i$ by developing complementary industries and leveraging its comparative advantages. Historically, norms and standards have evolved to restrain unilateral trade dominance; today, the World Trade Organization (WTO) plays a central role in regulating trade rules among nations.
@@ 4 Nash Bargaining Solutions for Reprisal Coefficients $\Lambda$
When additional factors are considered, the condition $\lambda_{ji}>\pi_{j}/\pi_{i}$ alone is insufficient for country $i$ to initiate a conflict against country $j$. Among these factors, trade deficits are particularly significant. This section explores bilateral bargaining solutions for $\lambda_{ji}$ when trade balance is also an objective.
@@ 4.1 Trade Balance
For any given period, a trade surplus for country $i$ with country $j$ implies an equivalent trade deficit for $j$ with $i$, and vice versa. A trade surplus creates job opportunities, reduces unemployment, and expands economies of scale. It also enhances creditworthiness by enabling debt repayment. Surpluses often result from comparative and competitive advantages in tradable goods and services. Opinions on trade deficits vary. While large deficits can undermine economic sustainability, in the short term, they may help avoid shortages and mitigate issues like inflation and poverty. The implications of trade deficits also depend on their impact on national security and how they are financed. No nation can completely ignore trade deficits or competitiveness; a one-sided emphasis is unsustainable.
There are two benchmark choices for $\lambda_{ji}$ regarding net trade balances: $g_{j}/g_{i}$ and $P_{ij}/P_{ji}$, as stated in Theorem 3. For country $i$ to have a trade surplus with country $j$ due to a change $\Delta P_{ij}$, $i$ must export more to $j$ than it imports, i.e., $g_{i}\Delta P_{ij}>g_{j}\Delta P_{ji}$. Thus, $\lambda_{ji}>g_{j}/g_{i}$ implies $i$’s trade surplus if $\Delta P_{ji}>0$. To sustain trade peace under $\Delta P_{ji}>0$, a large $\lambda_{ji}$ improves $i$’s trade balance but harms $\pi_{i}$ — excessive exports from $i$ to $j$ transfer power to $j$, reducing $\pi_{i}$. Therefore, $\lambda_{ji}$ should be neither too large nor too small. Conversely, $\lambda_{ji}<g_{j}/g_{i}$ also indicates $i$’s surplus if $\Delta P_{ji}<0$. In this case, both $i$ and $j$ lose exports, though $i$ loses less. Such outcomes can harm future production, so most countries avoid mutually damaging stalemates. At a cumulative level, $g_{i}P_{ij}$ and $g_{j}P_{ji}$ represent exports from $i$ to $j$ and from $j$ to $i$, respectively. Thus, $g_{j}/g_{i}=P_{ij}/P_{ji}$ results in zero trade surplus and deficit over the accumulation period. During this period, $g_{i}$, $g_{j}$, $P_{ij}$, and $P_{ji}$ all grow from zero, so $\Delta P_{ji}=P_{ji}-0$, $\Delta P_{ij}=P_{ij}-0$, and $\lambda_{ji}=\dfrac{\Delta P_{ij}}{\Delta P_{ji}}=\dfrac{P_{ij}}{P_{ji}}=\dfrac{g_{j}}{g_{i}}$.
@@
$\lambda_{ji}=g_{j}/g_{i}$ or $g_{j}/g_{i}=P_{ij}/P_{ji}$ maintains zero net trade deficit and surplus between countries $i$ and $j$ at the instantaneous or cumulative level, respectively.
To ease trade tensions, we consider two approaches for determining $\lambda_{ji}$: if both sides aim for higher competitiveness, then $\lambda_{ji}=\pi_{j}/\pi_{i}$ is the steady-state solution for any minor change in $P_{ji}$; if they seek balanced trade, then $\lambda_{ji}=g_{j}/g_{i}$ eliminates future trade deficits. In general, $\lambda_{ji}=\pi_{j}/\pi_{i}$ applies to long-term rivalry because $\pi$ reflects competitiveness across all countries and industries, whereas $\lambda_{ji}=g_{j}/g_{i}$ can immediately resolve trade deficits and alleviate related issues such as unemployment and national debt.
For example, according to Pozsar (2022), the 2018 China-USA trade war began in this manner: China had grown exponentially over four decades by exporting inexpensive goods overseas, while the USA lost much of its manufacturing base and accumulated massive national debt through wars on terror, significant tax cuts, and the 2007 subprime mortgage crisis. When China sought to build a global 5G network and produce cutting-edge computer chips, the USA attempted to block these moves and urged its allies to join the effort. In this case, the focal point was competitiveness rather than trade balance, suggesting a prolonged conflict where $\pi_{j}/\pi_{i}$ is the appropriate choice for $\lambda_{ji}$.
We examine a scenario in which a dominant superpower, country $i$, faces imminent challenges from an emerging economy, country $j$. According to the Matthew effect, it is likely that $\frac{\pi_{j}}{\pi_{i}}<\frac{g_{j}}{g_{i}}$. Based on Theorem 4, country $i$ cannot simultaneously achieve three objectives in its bilateral trade with country $j$: increasing its competitiveness $\pi_{i}$; maintaining a trade surplus; expanding exports to country $j$. This constraint creates what we call the Impossibility Trilemma. The trilemma implies that if country $i$ attempts to grow exports to country $j$ while increasing competitiveness, it cannot also sustain a trade surplus. This situation occurs when $\Delta P_{ij}<0$, causing both countries to lose comparative advantages.
@@
Given $\frac{\pi_{j}}{\pi_{i}}<\frac{g_{j}}{g_{i}}$, country $i$ cannot simultaneously achieve the three objectives in its bilateral trade with country $j$: increasing $\pi_{i}$; a trade surplus; positive export growth (i.e., $\Delta P_{ij}>0$).
Figure II illustrates this trilemma: at most two of the three policy objectives can be achieved by country $i$. For example, if country $i$ adopts a position along edge $b$, it can overturn comparative advantage and secure a trade surplus, but only by sacrificing competitiveness — unless it offsets this by absorbing competitiveness from other countries. A historical example is the early 2000s, when bilateral exports between China and the USA grew as a share of their respective production. According to the trilemma, the USA had to either lose competitiveness, incur trade deficits, or both.
Figure II: impossibility trilemma of $\pi_{i}\uparrow$, trade surplus, and $\Delta P_{ij}>0$
Country $i$’s interests in increasing competitiveness, maintaining a trade surplus, and achieving export growth are not perfectly aligned, and they do not completely contradict those of country $j$. In Figure II, country $i$’s optimal policy position lies somewhere within the triangle, while country $j$’s interests may lie elsewhere. A mutually beneficial outcome occurs when $\Delta P_{ij}>0$ and $\Delta P_{ji}>0$, as both countries gain from economies of scale and declining marginal production costs.
According to the proof of Theorem 4, country $j$ could, under certain conditions, achieve all three objectives simultaneously. While this may seem unfair to country $i$, it is conditional on the Matthew effect, which inherently favors $\pi_{i}$. Thus, Theorem 4 serves as a corrective measure to this bias.
@@ 4.2 Nash Bargaining Solutions
This subsection explores a compromise between trade balance and competitiveness. In international trade, some countries prioritize competitiveness, while others focus on improving trade balance — either by reducing deficits or increasing surpluses. These preferences influence the choice of $\lambda_{ji}$, which can maintain either competitiveness $\pi_{i}$ or trade balance unchanged at $\lambda_{ji}=\pi_{j}/\pi_{i}$ or $\lambda_{ji}=g_{j}/g_{i}$, respectively. If countries are indifferent between competitiveness and trade balance, a natural midpoint might seem like a solution: $\lambda_{ji}=\left(g_{j}/g_{i}+\pi_{j}/\pi_{i}\right)/2$ and $\lambda_{ij}=\left(g_{i}/g_{j}+\pi_{i}/\pi_{j}\right)/2$. However, this midpoint does not satisfy the identity condition $\lambda_{ji}\lambda_{ij}=1$.
When countries $i$ and $j$ have opposing priorities — one favoring competitiveness and the other trade balance — the Nash bargaining solution (1950) provides a cooperative approach that balances $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ while ensuring $\lambda_{ji}\lambda_{ij}=1$.
Define:
@@
which are either $\pi_{j}/\pi_{i}$ or $g_{j}/g_{i}$. Clearly,
@@
Without bargaining, according to Corollary 2 and Theorem 3, $\lambda_{ji}$ can secure $p$ or $q$, and $\lambda_{ij}$ can secure $1/p$ or $1/q$. These represent their respective status quo points if no agreement is reached.
Two scenarios arise when countries $i$ and $j$ have conflicting preferences. In the first scenario, country $j$ prefers $q$ over $p$, while $i$ prefers $1/p$ over $1/q$. However, $(q,1/p)$ is not a solution for $(\lambda_{ji},\lambda_{ij})$ since $q\times 1/p\neq 1$. By Eq. (15), $p$ and $1/q$ are their respective status quo points, obtained if one decides not to bargain with the other. In the second scenario, country $j$ prefers $p$ over $q$, while $i$ prefers $1/q$ over $1/p$. A preference over these alternatives defines a von Neumann-Morgenstern utility function for the bargaining problem.
Theorem 5 resolves these bargaining problems. The solution satisfies the identity $\lambda_{ji}\lambda_{ij}=1$ by design. It remains consistent across both scenarios and under any positive affine transformation of utility functions (a property of Nash bargaining). By taking the square root of $\pi_{j}/\pi_{i}$, the solution also mitigates the Matthew effect in $\pi$, bridging the gap between these $g_{j}/g_{i}$ and $\pi_{j}/\pi_{i}$.
@@
When countries $i$ and $j$ have opposing preferences on $p$ and $q$, the Nash bargaining solution is:
@@
Figure III illustrates the bargaining solution. In the first scenario (left plot), country $j$ moves from left to right along the curve $\lambda_{ij}=1/\lambda_{ji}$, while country $i$ moves from bottom to top along the same curve. They meet at the solution point $(\lambda^{*}_{ji},\lambda^{*}_{ij})$, which maximizes the area of the rectangle formed by this point and the status quo point $(p,1/q)$. The second scenario (right plot) shows a similar process where where $j$ moves from $q$ to $p$ and $i$ moves from $1/p$ to $1/q$ along the curve. Again, they meet at $(\lambda^{*}_{ji},\lambda_{ij}^{*})$, maximizing the area of the rectangle defined by this point and $(q,1/p)$.
Figure III: Nash bargaining solution for $\lambda_{ji}^{*}$ when $i$ and $j$ have opposite preferences
The solutions $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$ depend on the choice of status quo points. The effective bargaining set (or retaliation corridor) is $[p,q]$ for $\lambda_{ji}$ and $[1/q,1/p]$ for $\lambda_{ij}$, because country $j$ can secure $p$ (or $q$) and $i$ can secure $1/q$ (or $1/p$) in the first (or second) scenario. These corridors help prevent hostile overreactions and unintended consequences, such as third-party blowback. However, they can be adjusted if both players symmetrically strengthen or weaken their status quo positions.
If we scale the two status quo points by a positive number $c\leq\sqrt{q/p}$, the effective bargaining set for $\lambda_{ji}$ expands from $[p,q]$ to $[cp,q/c]$, and for $\lambda_{ij}$ from $[1/q,1/p]$ to $[c/q,1/(cp)]$. Despite this scaling, the bargaining solution remains unchanged (Corollary 4), while the new effective bargaining sets approach $(0,\infty)$ as $c\to 0$.
@@
For any positive $c\leq\sqrt{q/p}$, replacing $p$ with $cp$ and $q$ with $q/c$ does not change the bargaining solution: $\lambda_{ji}=\lambda_{ji}^{*}$ and $\lambda_{ij}=\lambda_{ij}^{*}$.
Additionally, as stated in Corollary 5, the ratios ${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ are also Nash bargaining solutions for $\lambda_{ji}$ when countries $i$ and $j$ directly compete for competitiveness and trade balance, respectively. The solution $\lambda_{ji}=\lambda_{ji}^{*}$ or $\pi_{j}/\pi_{i}$ — but not $g_{j}/g_{i}$ — appears endogenous to the derivative $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$, whose zero value provides a status quo point for the Nash solution. Nevertheless, the solution remains unchanged regardless of the derivative.
@@
Both ${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ are Nash bargaining solutions for $\lambda_{ji}$.
However, direct confrontation may be an inappropriate option when $g_{i}$ and $g_{j}$ differ significantly, according to Corollary 6. This aligns with China’s strategy with the USA from 1980 to 2010, when Deng Xiaoping advocated “keeping a low profile and taking no lead” (e.g., Yan, 2014). Historically, an even distribution of economic, political, and military capabilities between rival countries tends to increase the likelihood of war, whereas peace is best preserved when there is an imbalance of national capabilities (e.g., Organski, 1968). Therefore, Corollary 6 suggests using $\lambda_{ji}=\lambda_{ji}^{*}$ as the rule of thumb, unless their sizes of economies are comparable — in this case, $\frac{\pi_{j}}{\pi_{i}}$ and $\frac{g_{j}}{g_{i}}$ are also viable options for $\lambda_{ji}$.
@@
During globalization era when $\Delta P_{ji}>0$, country $j$ should strategically avoid homogeneous preferences with a significantly larger country $i$.